OpenStudy (anonymous):

Solve for n: 5n-1=2n+5 I'm having a math mental block please help!

OpenStudy (anonymous):

5n-1=2n+5 3n-1=5 3n=6 n=6/3=2

OpenStudy (anonymous):

5n-1=2n+5 1) subtract 2n from both sides then it would be 3n-1=5 2) add 1 to both sides then it would look like 3n=6 3)divide 3n into both sides then your answer would be n=2